Perform the following calculations to the correct number of significant figures.

Part A

(22.4856×2.55)+378.20

Part B

(86.5−22.060)÷0.0055

Part C

(582÷905.0)+5.35

Part D

[(2.68×106)÷53.589]+135.65
Answer #1

Few things we have to keep in mind before doing the problem:

  • Do all the operations before, then go for the result according to significant figure. If you consider significant figure in each and every step, then the final result will quite deviate from what it actually should be.
  • For multiplication and division result will have lowest significant figure among the numbers undergoing operation, for addition and subtraction result will have lowest decimal place among the numbers undergoing operation.
  • The significant figure of the result will be based on last operation performed.

A. (22.4856 x 2.55) + 378.20 = 57.33828 + 378.20 = 435.53828 = 435.54 (approximated to two decimal places)

B.  (86.5 − 22.060) / 0.0055 = 64.440 / 0.0055 = 11716.363 = 11716 (approximated to 5 significant figure)

C. (582 / 905.0) + 5.35 = 0.64 + 5.35 (here since we have no option so I took according to significant figure) = 5.99

D. [(2.68 × 106) / 53.589] + 135.65 = [284.08 / 53.589] + 135.65 = 5.3011 (approximated to 5 significant figure) + 135.65 = 140.9511 = 140.95 (approximated to 2 decimal points)